The marks obtained by 6 students in two class tests denoted by x and y, are as follows : 

x 6 5 8 10 4 9 

y 8 6 7 10 6 8 

Obtain the least square regression line of y on x.


1
Expert's answer
2021-07-19T18:56:04-0400

"y=mx+b"


"m=\frac{N\sum (xy)-\sum x\sum y}{N\sum x^2-(\sum x)^2}"


"b=\frac{\sum y-m\sum x}{N}"


Using calculator www.socscistatistics.com, we get:

"m=0.536, b=3.75"


"y=0.536x+3.75"